At Clucker's Chicken, a bucket of 4 pieces of dark meat and 5 pieces of white meat costs $7.05. A bucket of 3 pieces of dark meat and 8 pieces of white meat costs $8.90. Find the cost of a piece of dark meat and of a piece of white meat.

Answer by Boreal(15235)

4D+5W=7.05
3D+8W=8.90
-12D-15W=-21.15, multiplying by -3
12D+32W=35.60, multiplying by 4
17W=14.45, adding
W=$0.85
substitute into first to get D
4D+4.25=7.05
4D=2.80
D=$0.70
Check in second
2.10+6.80=8.90
white-$0.85; dark $0.70.